# Provable Ordering and Continuity in Vision-Language Pretraining for Generalizable Embodied Agents

### v1

@@ 5.6 Robustness Study under Linguistic Perturbations
In the EPIC-KITCHEN-100 dataset, textual annotations are often concise, such as “open cupboard”. In the default setting of LCBC, we employ similarly structured simple instructions. In this experiment, to validate the robustness of the representations our method learns in real-world scenarios, we introduce minor modifications to the language instructions. Specifically, we transform the original instruction “{action}” into two more conversational styles, i.e., “Please {action} for me.” and “Help me {action}.”. We then evaluate the imitation learning performance conditioned on these modified instructions in the Franka Kitchen environment. For comparison, we select LIV and DecisionNCE, which are also pre-trained on EPIC-KITCHEN-100. As shown in Table  2, The success rates of LIV and DecisionNCE dropped by 11.6% and 2.1%, respectively, whereas our method maintained a success rate comparable to that before the language perturbation. This result demonstrates the robustness of our learned representations, which generalize more effectively to real-world scenarios.
Table 2: Success rate fluctuation under linguistic perturbations on Franka Kitchen.
$\operatorname{Task}$ $\operatorname{LIV}$ $\operatorname{DecisionNCE}$ $\operatorname{AcTOL}$
Slide Cabinet $-29.0\pm 3.0$ $-4.5\pm 3.5$ $\mathbf{1.0}\pm 2.0$
Open Left Door $-3.5\pm 0.5$ $-1.5\pm 1.5$ $\mathbf{1.0}\pm 2.0$
Open Microwave $-4.5\pm 0.5$ $\mathbf{3.0}\pm 2.0$ $-2.5\pm 1.5$
Turn on stove $-8.5\pm 0.5$ $-6.5\pm 1.5$ $\mathbf{-0.5}\pm 1.5$
Switch on light $-12.5\pm 0.5$ $-1.0\pm 3.0$ $\mathbf{1.0}\pm 1.0$
$\operatorname{Average}$ $-11.6\pm 0.6$ $-2.1\pm 0.3$ $\mathbf{0.2}\pm 1.6$
6 Conclusion

### v3

@@ 5.3 Robustness Study under Linguistic Perturbations
In the EPIC-KITCHEN-100 dataset, textual annotations are often concise, such as ‘‘open cupboard’’. In the default setting of LCBC, we employ similarly structured simple instructions. In this experiment, to validate the robustness of the representations our method learns in real-world scenarios, we introduce several modifications to the language instructions. Specifically, we transform each original instruction into four conversational variants by varying lexical choices (e.g., verbs and nouns) and incorporating ChatGPT-4o [openai2024gpt4o] generated complex instructions. Details can be found in Appendix B.4. We then evaluate the imitation learning performance conditioned on these modified instructions in the Franka Kitchen environment. For comparison, we select LIV and DecisionNCE, which are also pre-trained on EPIC-KITCHEN-100. As shown in Figure 7, the success rates of LIV and DecisionNCE dropped by 11.9% and 2.7% on average, respectively, while our method maintained a success rate comparable to that before language perturbation. This result demonstrates the robustness of our learned representations, which generalize more effectively to real-world scenarios.
Figure 7: Success rate fluctuation across tasks in Franka Kitchen for different instruction variants.
5.4 Mitigating the Human-to-Robot Gap via Fine-Tuning
Table 4: Fine-tuning AcTOL encoders efficiently improves the success rate in Franka Kitchen.
Franka Kitchen Frozen Finetune
AcTOL 61.8 86.4
Although pretraining on human videos provides generalizable knowledge, bridging the human-to-robot domain gap remains a persistent challenge [zhou2024mitigating, li2025h2rhumantorobotdataaugmentation, nguyen2025robotic]. Since the AcTOL objectives capture the inherent temporal ordering of videos in an embodiment-agnostic manner, we can fine-tune the vision-language encoders with the same objectives used in pretraining during downstream behavior cloning. Notably, we find that fine-tuning with only a small number of robot demonstrations is sufficient to substantially mitigate the domain gap. We take 25 in-domain demonstrations (5 per task) in Franka Kitchen to fine-tune the pre-trained encoders using the AcTOL objectives. Then, as in Sec 5.1, we freeze the fine-tuned encoders and train policy networks on top using behavior cloning. We report the comparisons when using 15 demos for LCBC. As shown in Table 4, the success rate improvement demonstrates that the learned temporal inductive bias can be effectively adapted to the robot domain with limited supervision.
6 Related Work
Given the success of large-scale pre-training in the vision and language research communities [nips20-gpt3, nips23-llava], many studies have attempted to extend this paradigm to the field of robotics. Some work leverage massive robotic trajectory data [corr23-rtx] for pre-training, aiming to establish unified vision-langauge-action models [corl23-rt2, corr24-gr2, corr24-openvla, pi-zero, octo, DBLP:conf/nips/SzotMAHKT24, spatialvla]. However, collecting large amounts of high-quality robot trajectory data is extremely costly and time-consuming. Consequently, many studies have begun to explore the use of large-scale, readily available, out-of-domain human action video data to learn generalizable representations that can be transferred to robotic tasks [icra18-tcn, iclr23-vip, corl22-mvp, corl22-r3m, rss23-voltron, icml23-liv, nips23-vc1, DBLP:journals/corr/abs-2410-11758, rss24-mpi, icml24-decisionnce]. Among these, TCN [icra18-tcn], VIP [iclr23-vip], MVP [corl22-mvp], and VC-1 [nips23-vc1] focus solely on studying unimodal visual representations, limiting their performance when understanding language instructions is required. R3M [corl22-r3m] employs language and reward models to shape progressive visual representations, while Voltron [rss23-voltron] and MPI [rss24-mpi] model the transition from the current state to the goal state conditioned on language. However, during training, these approaches freeze the language encoder, using it only to aid in the training of visual representations. As a result, they do not effectively achieve multi-modal representation learning. LIV [icml23-liv] and DecisionNCE [icml24-decisionnce] have attempted to leverage CLIP [icml21-clip] to train embodied multi-modal representations. LIV treats language as the goal of video actions, aligning it with the final frame, while DecisionNCE aligns language with the transition from the initial to final frame. Both rely on a goal-reaching assumption, which can lead to suboptimal results in noisy real-world videos. In contrast, our approach avoids rigid assumptions by enforcing semantic alignment that follows the intrinsic temporal continuity of videos, leading to more robust and generalizable vision-language representations. This property also benefits methods like UVD [UVD], which rely on pretrained visual features to detect phase changes and decompose long-horizon tasks. Our method more reliably identifies action phases, enabling stronger progress rewards and improving suitability for such goal-conditioned downstream tasks.
7 Conclusion and Limitations